OpenStudy (anonymous):

Please help! Verify the identity: tanθcotθ-sin^2θ=cos^2θ

OpenStudy (noelgreco):

(tan x)(cot x) =1

OpenStudy (jdoe0001):

and \(\bf sin^2(\theta)+cos^2(\theta) =1\)

OpenStudy (anonymous):

so which on eis it? I need to show work too

OpenStudy (anonymous):

you need both to verify it

OpenStudy (anonymous):

first tan*cot =1 makes our equation 1-sin^20=cos^20

OpenStudy (anonymous):

then 1 = sin ^20+ cos^20

OpenStudy (anonymous):

which is true